# State of Health Estimation for Battery Modules with Parallel-Connected Cells Under Cell-to-Cell Variations

### 08 Dec 2023 v2

@@ III-C MI/CMI Estimation and Data Standardization
To have consistent estimation results, this paper uses the $k$-nearest neighbor ($k$NN) CMI estimator [30] to calculate both MI and CMI. Consider sample $\left\{\left(f_{i},g_{i},h_{i}\right)^{T}\right\}_{i=1}^{N}$ from RVs $\left(F,G,H\right)^{T}$ where $N$ is the number of sample points. For any sample point $\left(f_{i},g_{i},h_{i}\right)^{T}$, one can first define its $k$-nearest neighbor based on $l_{\infty}$-norm distance and denote the corresponding $l_{\infty}$-norm distance as $\frac{1}{2}d_{k,i}$. Note that $k$ is a tuning hyperparameter for the estimator and its proper value is related to the number of data [32]. Next, one can count the number of sample points within different windows:
@@ IV RELEVANCE VECTOR REGRESSION
Output: relevance vectors $\{\textbf{rv}_{j}\}_{j=1}^{N_{\text{rv}}}$, reciprocals of variances for offset and weights $\boldsymbol{\alpha_{\text{MP}}}$, noise variance $\beta_{\text{MP}}^{-1}$, trained/pruned $\boldsymbol{\Sigma}$, trained/pruned $\boldsymbol{\mu}$, a boolean variable $o$ about whether offset is used  Input: standardized data $\left\{\left(\boldsymbol{x}_{i},y_{i}\right)\right\}_{i=1}^{N}$, kernel function $K(\cdot,\cdot)$  Initialization: $\boldsymbol{\alpha}$ with $\alpha_{i}\leftarrow\frac{1}{\left(N+1\right)^{2}},i=0,...,N$, $\boldsymbol{\alpha}^{\text{old}}$ with $\alpha^{\text{old}}_{i}\leftarrow\alpha_{i}$, $\beta^{-1}\leftarrow\left(0.1\cdot\text{std}\left(\{y_{i}\}_{i=1}^{N}\right)\right)^{2}$, a set $\mathcal{E}\leftarrow\left\{e_{0},\boldsymbol{x}_{1},...,\boldsymbol{x}_{N}\right\}$ ($e_{0}$ serves as a placeholder for offset and can be initialized as anything), maximum number of iterations $N_{\text{iter}}$, threshold $\alpha_{\text{th}}$ (a very large number), tolerance tol, a very small number $\varepsilon$, $\boldsymbol{\Phi}\leftarrow\begin{bmatrix}\boldsymbol{\phi}\left(\boldsymbol{x}_{1}\right)&...&\boldsymbol{\phi}\left(\boldsymbol{x}_{N}\right)\end{bmatrix}^{T}$ where $\boldsymbol{\phi}\left(\boldsymbol{x}_{i}\right)\leftarrow\begin{bmatrix}1&K\left(\boldsymbol{x}_{i},\boldsymbol{x}_{1}\right)&...&K\left(\boldsymbol{x}_{i},\boldsymbol{x}_{N}\right)\end{bmatrix},\forall i=1,...,N$  for $k=1,...,N_{\text{iter}}$ do
@@
With the optimal set of features found, the next step is to build a module level SOH estimation model. This section corresponds to Step 4 of Fig. 2 where the proposed algorithm trains a sparse probabilistic model and estimates module level SOH as distributions with point estimates and three-sigma credible intervals.
Relevance vector regression (RVR) is a sparse regression technique and the Bayesian counterpart to non-Bayesian support vector regression (SVR) [39]. RVR has the following advantages. First, trained models can be updated based on future data using sequential Bayesian learning [39]. Second, when making predictions, instead of seeking to have accurately trained model parameters, the RVR integrates over all these unknown parameters, which improves predictive performance and increases robustness [39]. Third, the RVR automatically prefers simple but sufficiently accurate models over complicated models to explain the data [40]. Fourth, kernels in RVR do not need to satisfy Mercer’s conditions, allowing higher design freedom [41].
@@ IV-A Regression Algorithm
Assume independent and identically distributed (I.I.D.) data $\left\{\left(\boldsymbol{x}_{i},y_{i}\right)\right\}_{i=1}^{N}$ where $\boldsymbol{x}\in\mathbb{R}^{N_{f}}$ are features, $N_{f}$ is the number of features, $y\in\mathbb{R}$ is the output (i.e., SOH), and $N$ is the number of data. Then, an output $y$ and an input $\boldsymbol{x}$ can be related by:
@@
The next step is to choose a likelihood distribution for the given data [42]. Based on the assumptions of I.I.D. sample points and white noises in the process, the likelihood for RVR is:
@@ V BATTERY DATASETS
Cell Level SOH Range: 100% - 88%,
@@
Cell Experiment: Detailed descriptions are given in [46].
Downloading Website: [47]
@@
Module Level SOH Range: 100% - 86%
Module Experiment: The entire battery pack containing 96 modules in series is cycled under the charging profiles consisting of two consecutive constant-current regimes with 86A and 64.5A pack-level currents, respectively, and one subsequent constant-voltage regime. Only the first constant-current segment is used for this study, as this segment contains all the IC and DV features for this dataset.
The proposed method in Fig. 2 will be demonstrated and validated using experimental datasets. First, the public Oxford lithium cobalt oxide (LCO) cell dataset [46] is used to illustrate related concepts and interpret results from the proposed method. Then, a proprietary lithium nickel-manganese-cobalt oxide 622 (NMC622) module dataset is used to evaluate the performance of the proposed method for module level SOH estimation under cell-to-cell variations. Table II summarizes the key attributes of these two datasets, while Fig. 4 shows their example IC and DV curves and related features. Note that, for ease of discussion, the acronyms in Table I and indices labeled in Fig. 4 will be used together to refer to different features. For example, the left IC peak height is denoted as IC PH 1.
@@
Figure 5: Feature Relevance, Redundancy, and Complementarity for Cell Level SOH Estimation under Different C Rates
@@ VI INTERPRETATION OF THE PROPOSED METHOD USING CELL DATA
The Oxford cell level dataset is split into 80% training and 20% testing sets. The training data are used for performing the feature selection and building the RVR-based SOH estimation model, while testing data are used to evaluate the performance.
@@
The black color in Fig. 4(b) represents two features with redundancy above the threshold $\tilde{I}_{\text{th}}$. The darker the color is, the higher the redundancy is. From Fig. 4(b), one can observe that:
  • IC peak/valley heights and DV valley/peak heights are completely redundant, because $\text{IC}=1/\text{DV}$.
@@
  • C Rate is completely redundant to IC PL 2, IC VH 2, IC VL 1, and DV PH 2. Note that this observation is only applicable to this dataset.
@@
TABLE III: Feature Selection Results for Cell Level SOH Estimation of LCO Cells under Different C Rates
@@
(On Values of Normalized CMI) Note that, because of the estimation inaccuracy of the used CMI estimator, the normalized CMI could occasionally become a little above 1 as shown in Fig. 4(c).
@@
These top two ranked features are used to build a cell level SOH estimation model. Based on Table IV, the proposed model has a low computational complexity (small number of relevance vectors, $\textbf{rv}_{j}$) and high estimation accuracy (low test root-mean-square error, RMSE). Fig. 6 shows the trained model under different C rates and the three-sigma (99.7%) credible intervals for different SOHs. Table IV also summarizes the average three-sigma values.
Figure 6: Cell Level SOH Estimation Model Using Two Features
(a) Estimation RMSE
(b) Average Three-Sigma Values
(c) Number of Relevance Vectors
Figure 7: Module Level SOH Estimation Performance under Different Number of Features Used
TABLE IV: Cell Level SOH Estimation Results
@@
VII MODULE LEVEL SOH ESTIMATION RESULTS UNDER CELL-TO-CELL VARIATIONS
This section presents the results of module level SOH estimation in the presence of cell-to-cell variations, using the experimental NMC622 module data to demonstrate real-world performance.
@@ VII MODULE LEVEL SOH ESTIMATION RESULTS UNDER CELL-TO-CELL VARIATIONS
The proposed Algorithm 1 finds the optimal set of features for module level SOH estimation in the presence of cell-to-cell variations. Table V summarizes the results. Note that, with different conditions from Section VI, the proposed feature selection algorithm finds another different optimal set of ranked features. IC PA 1 is the top-ranked feature, which is not only related to SOH but also the most insensitive to cell-to-cell variations.
TABLE V: Feature Selection Results for Module Level SOH Estimation in the Presence of Cell-to-Cell Variations
@@
TABLE VI: Module Level SOH Estimation Results for Modules with Cell-to-Cell Variations under Different Number of Features Used
@@
(with the feature used in cell level SOH estimation [13, 14, 15, 18])
@@
The final SOH estimation model is trained using all the training data, and Table VI summarizes the final testing RMSE, number of relevance vectors, and average three-sigma values when different numbers of features are used. Note that, based on Table VI, even the one-feature SOH estimation model using IC PA 1 outperforms the conventional IC peak-based SOH estimation model originally developed for the cell level, in terms of estimation accuracy, estimation confidence, and computational efficiency.
@@
Fig. 7 shows how the RMSE, number of relevance vectors, and average three-sigma values vary with the number of features used in the SOH estimation model for cross-validation and testing. Fig. 8 shows the absolute estimation error distribution for all the testing data when only two features are used. Based on Table VI, Fig. 7, and Fig. 8, several important conclusions could be made:
  • Leveraging properly selected multiple features indeed improves module level SOH estimation accuracy and reduces estimation uncertainty.
  • Better performance for module level SOH estimation can be achieved at the expense of higher model complexity. Using only two or three features provides a good trade-off between estimation performance and model complexity.
  • Module level SOH can indeed be estimated with high accuracy and confidence in the presence of cell-to-cell variations. Note that the proposed method does not use any information about cell-to-cell variations explicitly.
@@ VIII ONBOARD IMPLEMENTATION OF THE PROPOSED METHOD
The proposed method in Fig. 2 has both offboard and onboard computations involved. The proposed feature selection algorithm in Step 3 of Fig. 2 is performed completely offboard to find the optimal set of features for module level SOH estimation. The onboard computation involves: (1) extracting the values of these selected features from module level IC and DV curves and (2) estimating SOH according to the RVR model developed in Step 4 of Fig. 2.
@@ IX CONCLUSIONS
This paper proposes a novel method and demonstrates the feasibility of estimating module level SOH with high accuracy and confidence in the presence of cell-to-cell variations. First, a novel information theory-based feature selection algorithm is proposed to find an optimal set of features for SOH estimation under cell-to-cell variations by optimizing feature relevance, redundancy, and complementarity. The optimal feature set found is independent of subsequent learning algorithms. Second, a novel relevance vector regression (RVR)-based SOH estimation model is proposed. Model sparsity, high estimation accuracy, and good estimation confidence are demonstrated. Applied to a large dataset, the proposed algorithm achieves module level SOH estimation with 0.53% RMSE and 1.56% average three-sigma value when only two features are used. With more ranked features used in the estimation model, the accuracy can be improved to 0.48% RMSE and 1.42% average three-sigma value. Thus, compared to the estimation results (1.3% RMSE and 3.9% average three-sigma value) obtained by directly applying IC peak-based methods originally developed for the cell level, the proposed method provides significant improvements in estimation performance. As the optimization and training processes are performed offboard, the developed SOH model can be implemented for onboard execution with low memory and computational requirements.
Future work on the following topics will be pursued to extend and generalize the results reported in this paper. First, making the feature selection algorithm more robust under different operating conditions will be investigated. Second, the effects of measurement noise, data acquisition systems, and vehicle chronometrics will be investigated. Third, different approaches to estimating cell-to-cell variations within the module, given only module level measurements, will be explored.

### 04 Apr 2024 v4

@@ III-C MI/CMI Estimation and Data Standardization
To have consistent estimation results, this paper uses the $k$-nearest neighbor ($k$NN) CMI estimator [33] to calculate both MI and CMI. Consider sample $\left\{\left(f_{i},g_{i},h_{i}\right)^{T}\right\}_{i=1}^{N}$ from RVs $\left(F,G,H\right)^{T}$ where $N$ is the number of sample points and $i$ is an index. For any sample point $\left(f_{i},g_{i},h_{i}\right)^{T}$, one can first define its $k$-nearest neighbor based on $l_{\infty}$-norm distance and denote the corresponding $l_{\infty}$-norm distance as $\frac{1}{2}d_{k,i}$. Note that $k$ is a tuning hyperparameter for the estimator and its proper value is related to the number of data [35]. Next, one can count the number of sample points within different windows:
@@ IV RELEVANCE VECTOR REGRESSION
Output: relevance vectors $\{\textbf{rv}_{j}\}_{j=1}^{N_{\text{rv}}}$, reciprocals of variances for offset and weights $\boldsymbol{\alpha_{\text{MP}}}$, noise variance $\beta_{\text{MP}}^{-1}$, trained/pruned $\boldsymbol{\Sigma}$, trained/pruned $\boldsymbol{\mu}$, a boolean variable $o$ about whether offset is used  Input: standardized data $\left\{\left(\boldsymbol{x}_{i},y_{i}\right)\right\}_{i=1}^{N}$, kernel function $K(\cdot,\cdot)$  Initialization: $\boldsymbol{\alpha}$ with $\alpha_{i}\leftarrow\frac{1}{\left(N+1\right)^{2}},i=0,...,N$, $\boldsymbol{\alpha}^{\text{old}}$ with $\alpha^{\text{old}}_{i}\leftarrow\alpha_{i}$, $\beta^{-1}\leftarrow\left(0.1\cdot\text{std}\left(\{y_{i}\}_{i=1}^{N}\right)\right)^{2}$, a set $\mathcal{E}\leftarrow\left\{e_{0},\boldsymbol{x}_{1},...,\boldsymbol{x}_{N}\right\}$ ($e_{0}$ serves as a placeholder for offset and can be initialized as anything), maximum number of iterations $N_{\text{iter}}$, threshold $\alpha_{\text{th}}$ (a large number), tolerance tol, a small number $\varepsilon$, $\boldsymbol{y}=\begin{bmatrix}y_{1}&...&y_{N}\end{bmatrix}^{T}$, $\boldsymbol{\Phi}\leftarrow\begin{bmatrix}\boldsymbol{\phi}\left(\boldsymbol{x}_{1}\right)&...&\boldsymbol{\phi}\left(\boldsymbol{x}_{N}\right)\end{bmatrix}^{T}$ where $\boldsymbol{\phi}\left(\boldsymbol{x}_{i}\right)\leftarrow\begin{bmatrix}1&K\left(\boldsymbol{x}_{i},\boldsymbol{x}_{1}\right)&...&K\left(\boldsymbol{x}_{i},\boldsymbol{x}_{N}\right)\end{bmatrix},\forall i=1,...,N$  for $k=1,...,N_{\text{iter}}$ do
@@
With the optimal set of features found, the next step is to build a module-level SOH estimation model. This section corresponds to Step 4 of Fig. 2 where the proposed algorithm trains a sparse probabilistic model and estimates module-level SOH as distributions with point estimates and three-sigma credible intervals.
Relevance vector regression (RVR) is a sparse regression technique and the Bayesian counterpart to non-Bayesian support vector regression [42]. RVR has the following advantages. First, trained models can be updated based on future data using sequential Bayesian learning [42]. Second, when making predictions, instead of seeking to have accurately trained model parameters, the RVR integrates over all these unknown parameters, which improves predictive performance and increases robustness [42]. Third, the RVR automatically prefers simple but sufficiently accurate models over complicated models to explain the data [43]. Fourth, kernels in RVR do not need to satisfy Mercer’s conditions, allowing higher design freedom [44].
@@ IV-A Regression Algorithm
Assume independent and identically distributed (I.I.D.) data $\left\{\left(\boldsymbol{x}_{i},y_{i}\right)\right\}_{i=1}^{N}$ where $\boldsymbol{x}\in\mathbb{R}^{N_{f}}$ are features, $N_{f}$ is the number of features, $y\in\mathbb{R}$ is the output (i.e., SOH), $i$ is an index, and $N$ is the number of data. Then, an output $y$ and an input $\boldsymbol{x}$ can be related by:
@@
The next step is to choose a likelihood distribution for the given data [45]. Based on the assumptions of I.I.D. sample points and white Gaussian noises in the process, the likelihood for RVR is:
@@ V BATTERY DATASETS
The proposed method in Fig. 2 will be demonstrated and validated using experimental datasets. First, the public Oxford lithium cobalt oxide (LCO) cell dataset [49] (downloading website: [50]) is used to illustrate related concepts and interpret results from the proposed method. The detailed description can be found in [49]. Then, a proprietary lithium nickel-manganese-cobalt oxide 622 (NMC622) module dataset is used to evaluate the performance of the proposed method for module-level SOH estimation under cell-to-cell variations. The existence of cell-to-cell variation within modules and their impacts on IC curves are illustrated in Fig. 1. The entire battery pack containing 96 modules in series is cycled under the charging profiles consisting of two consecutive constant-current regimes with 86A and 64.5A pack-level currents, respectively, and one subsequent constant-voltage regime. Only the first constant-current segment is used for this study, as this segment contains all the IC and DV features for this dataset. The measurements are collected at a 1Hz rate, but they are downsampled so that the actual data used by the proposed method have a 0.033Hz sampling rate, i.e., 1 measurement per 30 seconds. Passive balancing is used for serially connected modules. The balancing current is very small compared to the module-level current, thus its impact on IC and DV curves is very small. Parallelly connected cells have self-balancing effects and do not involve active or passive cell balancing.
@@
Cell-Level SOH Range: 100% - 88%,
@@
Module-Level SOH Range: 100% - 86%
@@
Figure 5: Feature Relevance, Redundancy, and Complementarity for Cell-Level SOH Estimation under Different C Rates
Table II summarizes the key attributes of these two datasets, while Fig. 4 shows their example IC and DV curves and related features. Note that, for ease of discussion, the acronyms in Table I and indices labeled in Fig. 4 will be used together to refer to different features. For example, the left IC peak height in Fig. 3(a) is denoted as IC PH 1.
@@ VI INTERPRETATION OF THE PROPOSED METHOD USING CELL DATA
The cell-level SOH estimation is discussed first, because the cell dataset contains a lot more features and charging conditions than the proprietary module dataset, making the cell dataset more suitable to demonstrate the full capability of the proposed feature selection algorithm. The Oxford cell dataset is split into 80% training and 20% testing sets. The training data are used for performing the feature selection and building the RVR-based SOH estimation model, while testing data are used to evaluate the performance.
@@
TABLE III: Feature Selection Results for Cell-Level SOH Estimation of LCO Cells under Different C Rates
@@
(On Values of Normalized CMI) Note that, because of the estimation inaccuracy of the used CMI estimator, the normalized CMI could occasionally become a little above 1, as shown in Fig. 4(c).
@@
These top two ranked features are used to build a cell-level SOH estimation model. Based on Table IV, the proposed model has a low computational complexity (small number of relevance vectors, $\textbf{rv}_{j}$) and high estimation accuracy (low test root-mean-square error, RMSE). Fig. 6 shows the trained model under different C rates and the three-sigma (99.7%) credible intervals for different SOHs. Table IV also summarizes the average three-sigma values.
TABLE IV: Cell-Level SOH Estimation Results
@@
Figure 6: Cell-Level SOH Estimation Model Using Two Features
(a) Estimation RMSE
(b) Average Three-Sigma Values
(c) Number of Relevance Vectors
Figure 7: Module-Level SOH Estimation Performance under Different Number of Features Used
VII MODULE-LEVEL SOH ESTIMATION RESULTS UNDER CELL-TO-CELL VARIATIONS
This section presents the results of module-level SOH estimation in the presence of cell-to-cell variations, using the experimental NMC622 module data to demonstrate real-world performance.
@@ VII MODULE-LEVEL SOH ESTIMATION RESULTS UNDER CELL-TO-CELL VARIATIONS
The proposed Algorithm 1 finds the optimal set of features for module-level SOH estimation in the presence of cell-to-cell variations. Note that the feature selection algorithm is the same for both cell- and module-level SOH estimations. This is the beauty of this proposed method — one unified method for two tasks. Nonetheless, the selected optimal sets of features are different for cell- and module-level SOH estimations, as the situations and conditions change.
Table V summarizes the module-level feature selection results. IC PA 1 is the top-ranked feature, which is not only related to SOH but also the most insensitive to cell-to-cell variations. Other features are selected, according to the trade-off they provide among relevance, redundancy, and complementarity, instead of how insensitive they are to cell-to-cell variations. Referring back to the issue shown in Fig. 1, even though IC PH 1 is significantly distorted by cell-to-cell variations, the proposed method finds that IC PH 1 provides the best trade-off among relevance, redundancy, and complementarity, once IC PA 1 and IC PA 2 are selected. Thus, instead of using IC PH 1 as a sole feature for SOH estimation, the proposed algorithm recommends IC PH 1 to be used together with IC PA 1 and IC PA 2 to provide optimal estimation performance for this dataset.
TABLE V: Feature Selection Results for Module-Level SOH Estimation in the Presence of Cell-to-Cell Variations
@@
TABLE VI: Module-Level SOH Estimation Results for Modules with Cell-to-Cell Variations under Different Number of Features Used
@@
(with the feature used in cell-level SOH estimation [13, 14, 15, 18])
@@
The final SOH estimation model is trained using all the training data, and Table VI summarizes the final testing RMSE, number of relevance vectors, and average three-sigma values when different numbers of features are used. Based on Table VI, even the SOH estimation model using only IC PA 1 (selected by the proposed feature selection algorithm) significantly outperforms the conventional SOH estimation model using only IC PH 1 (selected based on literature without feature selection), in terms of estimation accuracy, estimation confidence, and computational efficiency. From another perspective, these results also exemplify the significant extent to which cell-to-cell variations can impact SOH estimation. IC PA 1 (the most robust feature against cell-to-cell variation for this dataset) can achieve roughly 1/2 of both the test RMSE and average three-sigma value that IC PH 1 (commonly used for cell-level SOH estimation but is sensitive to cell-to-cell variations at the module level) achieves.
To reflect the trade-off between estimation performance and model complexity, Fig. 7 shows how the RMSE, number of relevance vectors, and average three-sigma values vary with the number of features used in the SOH estimation model for cross-validation and testing. Fig. 8 shows the absolute estimation error distribution for all the testing data when only two features are used. Based on Table VI, Fig. 7, and Fig. 8, several important conclusions could be made:
  • Leveraging properly selected multiple features indeed improves module-level SOH estimation accuracy and reduces estimation uncertainty.
  • Better performance for module-level SOH estimation can be achieved at the expense of higher model complexity. Using only two or three features provides a good trade-off between estimation performance and model complexity.
  • Module-level SOH can indeed be estimated with high accuracy and confidence in the presence of cell-to-cell variations. Note that the proposed method does not use any information about cell-to-cell variations explicitly.
@@
Remark 5.
For the proposed SOH estimation method, charging ranges do not necessarily have to be full but have to be long enough to include the selected features. Since the SOH estimation, unlike the state of charge estimation, does not need to be performed continuously for all operations, a condition can be included in the algorithm so that SOH estimation is only performed when there are sufficient data points and bypassed when charging ranges are short.
@@ VIII ONBOARD IMPLEMENTATION OF THE PROPOSED METHOD
The proposed method in Fig. 2 has both offboard and onboard computations involved. The proposed feature selection algorithm in Step 3 of Fig. 2 is performed completely offboard to find the optimal set of features for module-level SOH estimation. The onboard computation involves: (1) extracting the values of these selected features from module-level IC and DV curves and (2) estimating SOH according to the RVR model developed in Step 4 of Fig. 2.
@@ IX CONCLUSIONS
This paper proposes a novel method and demonstrates the feasibility of estimating module-level SOH with high accuracy and confidence in the presence of cell-to-cell variations. First, an information theory-based feature selection algorithm is proposed to find an optimal set of features for SOH estimation under cell-to-cell variations by optimizing feature relevance, redundancy, and complementarity. The optimal feature set found is independent of subsequent learning algorithms. Second, a relevance vector regression (RVR)-based SOH estimation model is proposed. Model sparsity, high estimation accuracy, and good estimation confidence are demonstrated. Applied to a large dataset, the proposed algorithm achieves module-level SOH estimation with 0.53% RMSE and 1.56% average three-sigma value when only two features are used. With more ranked features used in the estimation model, the accuracy can be improved to 0.48% RMSE and 1.42% average three-sigma value. Thus, compared to the estimation results (1.3% RMSE and 3.9% average three-sigma value) obtained by directly applying IC peak-based methods originally developed for the cell level, the proposed method provides significant improvements in estimation performance. As the optimization and training processes are performed offboard, the developed SOH model can be implemented for onboard execution with low memory and computational requirements.
Results in Sections VI and VII indicate that the same proposed method in Fig. 2 has good generalizability to work well for both cell-level and module-level SOH estimation problems. Different optimal sets of features will be found by the proposed method for cell-level and module-level SOH estimation problems to give good estimation performance.
Future work on the following topics will be pursued to extend and generalize the results reported in this paper. First, making the feature selection algorithm more robust under different operating conditions will be investigated. For example, two experimental datasets used in Sections VI and VII verify the proposed method in the situation where all the originally available features will not disappear throughout the entire degradation process. Future research could be conducted to see how the proposed method performs when some originally available features disappear at some degradation status. Second, how to extend the proposed method to different charging ranges, especially those that do not contain features recommended by the feature selection algorithm can be investigated in the future. Third, the effects of measurement noise, data acquisition systems, and vehicle chronometrics on the proposed method will be investigated. Fourth, different approaches to estimating cell-to-cell variations within the module, given only module-level measurements, will be explored. Fifth, whether the proposed framework could be applied to estimate the internal resistances of cells and modules for battery power-related performance could also be investigated in the future.